OpenStudy (anonymous):

Kelly cannot remember the formula for the sum of a finite arithmetic sequence, so she is going to try to derive it. Her problem is to find the sum of the numbers 1 to 500. What should she do first?

OpenStudy (anonymous):

add forwards and backwards \(S=1+2+3+...+498+499+500\) \(S=500+499+498+...+3+2+1\) add together to get \[2S=501+501+501+...+501+501+501=500\times 501\]

OpenStudy (anonymous):

Thanks!!